OpenStudy (anonymous):

N/F = n/f If i have N and n is proportional to C and c, can i say C/F = c/f ?

OpenStudy (phi):

if \[ \frac{N}{F} = \frac{n}{f} \text{ then } \\ \frac{N}{n}= \frac{F}{f} \] you show this by cross multiplying to get N f = F n and then divide both sides by f n to get N/n = F/f if you also have \[ \frac{N}{C} = \frac{n}{c} \] then you have \[ \frac{N}{n} = \frac{C}{c} \] Because \[ \frac{N}{n}= \frac{F}{f}\text { and } \frac{N}{n} = \frac{C}{c} \] you can say \[ \frac{F}{f}= \frac{C}{c} \] which can be re-arranged. cross multiply \[ F c = C f \] divide both sides by F f to get \[ \frac{F c}{F f} = \frac{C f}{F f} \\ \frac{c}{f}= \frac{C}{F}\]